OpenStudy (promisespromises):

Find f ' (x) if f is defined by f(x) = xsin(1/x) and if x cannot = 0 and f(0) =0

OpenStudy (bittuaryan):

You mean: \[\frac{ d }{ dx }xsin \left( \frac{ 1 }{ x } \right)\]

OpenStudy (bittuaryan):

=(-2/x)cos(1/x)+sin(1/x)

OpenStudy (promisespromises):

I wrote it exactly as it is on my review sheet

OpenStudy (promisespromises):

I'm pretty sure I should use the product rule. The \[x \neq 0\] threw me off though.